Determine pH of a weak base that has Kb = 3.78 x10-18
1
Expert's answer
2020-04-27T01:51:17-0400

Solution:

The equation representing the ionization of a weak base (B):

B(aq) + H2O = BH+(aq) + OH-(aq)



where Co = Co(B) - initial concentration of a weak base solution.

[OH-] = [BH+] = x

[B] = Co(B) - x

In the problem statement, the initial concentration of a base solution is not provided (although it should be), therefore:

Kb = [x] * [x] / [Co(B) - x] = 3.78*10-18

Knowing the values of Co(B), we substitute it in the quadratic equation:

x2 + x*3.78*10-18 - Co(B)*3.78*10-18 = 0.

Solving the quadratic equation, we find the values of x.

Then,

x= [OH-]

pOH = -log[OH-] = -log(x).

pH = 14 - pOH

pH = 14 + log(x).